OpenStudy (anonymous):

The length of a rectangular floor is 2 feet more than its width. The area of the floor is 168 square feet. Kim wants to use a rug in the middle of the room and leave a 2-foot border of the floor visible on all sides. What should the width (the shorter side) of the rug be?

OpenStudy (anonymous):

@papiblow_17 Hey yeah, nice to Meet you !! Welcome to Open Study !! 1. Find the dimensions of the room .... x(x+2) = 168 x^2 + 2x - 168 = 0 (x-12)(x+14) = 0 continue yourself.. x= ? Solve and Tag Anyone who Knows, and Verify your answer !! Have A Nice Day Cheers !!